sadecummings:

Which number can each term of the equation be multiplied by to eliminate the fractions before solving? -3/4m-1/2=2+1/4m

AZ:

is this the equation \(-\dfrac{3}{4} m - \dfrac{1}{2} = 2+ \dfrac{1}{4}m\)

AZ:

look at the denominator you have a 4 and a 2 what do you think you need to multiply on both sides to get rid of the 4 in the denominator?